A planet orbits a star, in a year of length 2.23 x 107 s, in a nearly circular orbit of radius 2.89 x 1011 m. With respect to the star, determine (a) the angular speed of the planet, (b) the tangential speed of the planet, and (c) the magnitude of the planet's centripetal acceleration.
Answer #1

a)

angular velocity

w = 2 pi / T = 2 pi / ( 2.23* 10^7)

w = 2.816* 10^-7 rad/s

=====

b)

v = r w = 2.89* 10^11* 2.816* 10^-7 = 81386.55 m/s

=====

c)

a = r w^2 = 0.0229 rad/s^2

=====

Comment before rate in case any doubt, will reply for sure.. goodluck